OpenStudy (anonymous):

HELP a = 16 b = 182 c = 0 find the coordinates of the vertex.

OpenStudy (anonymous):

i have tried so many times and get weird awnsers

OpenStudy (anonymous):

what do those numbers suppose to represent?

OpenStudy (anonymous):

16x^ 2 + 182x

OpenStudy (anonymous):

16x^2 + 182x

OpenStudy (anonymous):

then i think you put it like x = -b/2a

OpenStudy (anonymous):

oh... those are the coefficients for a parabola in standard form: \(\large y=ax^2+bx+c \)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

you are correct... the x-coordinate of the vertex is \(\large \frac{-b}{2a} \)

OpenStudy (anonymous):

then i get -5.6875

OpenStudy (anonymous):

yes, but leave it in fraction form... reduced....

OpenStudy (anonymous):

oh what difference does that make?

OpenStudy (anonymous):

many times the exact value can only be written in fraction form...

OpenStudy (anonymous):

well i think it might end up better than -1126.125 right?

OpenStudy (anonymous):

or in the end it will come out the same?

OpenStudy (anonymous):

idk what you mean ^^^

OpenStudy (anonymous):

when i plug -5.6875 into 16x^2 + 182x it makes -1126

OpenStudy (anonymous):

.125

OpenStudy (anonymous):

oh... yeah ok... if you're teacher doesn't mind it, then i guess it's ok...

OpenStudy (anonymous):

so that is the answer? the order pair is (-5.6875, -1126.125)

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

because the question is about how high water can be shot

OpenStudy (anonymous):

from a jet

OpenStudy (anonymous):

How High Will the Water Reach? 10. Identify the a, b, and c values in your equation. (3 points) a = 16 b = 182 c = 0 11. What is the formula for locating the x-coordinate of the vertex? (2 points) 16x2 + 182x 12. Use the graphing technology of your choice to graph the above equation and find the vertex value. Set rounding to 5 places. Save a copy of this graph and paste or attach it to this Worksheet. (5 points) 13. What are the coordinates of the vertex? (5 points) (5.6875, 1126.125) 14. How high does this mean the water from your fountain reaches into the air? (5 points) It reaches 1126.125 feet in the air. 15. Does this match the height indicated by your research? (1 point) • Yes • No 16. If it is different, what do you think may be a possible reason why the two numbers differ? (3 points)

OpenStudy (anonymous):

for that y-value, i got something else... double check your answer.

OpenStudy (anonymous):

and how come in your answer you changed the x-coordinate from negative to positive?

OpenStudy (anonymous):

i just plug in the x right? i got a positive by multiplying -91 and -1035.125

OpenStudy (anonymous):

are you sure that 16 is supposed to be positive? is the coefficient of x^2 supposed to be 16?

OpenStudy (anonymous):

oh its - 16

OpenStudy (anonymous):

so the only difference now is that it is negative.